Transition from morning work (*:responsible person)
- Current status as of 1/22: IR and Green reached X end PD.
1/22 Surface inspection for PR mirrors. Cleaning will be done after 1/31.
- 1/22 or 23 Surface inspection for SR mirrors.
- 1/23 Cleaning HR of SR2.
- 1/23 BS will be aligned to Y-end PD by IR.
- 1/24-25 ITMX will be aligned to REFL by IR.
- 1/24-25 ITMY will be aligned to REFL by IR.
- 1/26 Beam position of IR at SR3 will be confirmed how far from the center.
- 1/28-29 SR3 will be aligned to SR2 center by IR.
- 1/30 Green beam will be injected from SR2 AR surface to Y end PD.
- 1/31-2/1 Removing bellows at PR and SR area.
- 1/31-2/1 Cleaning HR of PR3 and AR of SR2
2/4-8 Alignment of SR2 -> SRM by IR, SRM -> OMMT by IR -> OMMT -> OSTM by IR, OSTM -> OMC by IR
2/4-8 Alignment of ETMY -> POS by Green
- 2/12-15 Closing door for central area, pumping down
- needs drift monitor for VIS
Report of work (*:responsible person)
- BS: Two cables were not connected well, so actuation was not correct when measuring TF day before yesterday.
- TF was measured yesterday but still TF showed something different from the last correct measurement.
- FR moved 800um.
- Anyway alignment can be proceeded probably this afternoon.
- CRY: we found disconnected cables at ITMX.
- OMC: OMMT was aligned, and it was found that OSEM was touched somewhere.
- PEM: Ondotori is still not working because of software issue.
Work plan (*:responsible person)
- COMM: we will check PR3 first, then SR3.
- Access to PR3 is quite difficult, gate valve needs to be removed or optical table needs to be fix.
- we will decide whether we need clean the mirror surface or not.
- Once mirror inspection and cleaning (if necessary), then will start alignment for REFL using reflected beam from ITMX.
- COMM: PDA will be placed EYA.
- SR3, SR2: height is OK. FR will be tuned.
- SR3 surface will be inspected.